Compound X has a molecular formula C8H10 and gives the 1H NMR spectrum below. What is the structure of X?

11 10 11 10 98 76 54 3 2 1 0 ppm 1H NMR
Answer #1

in the given molecular formula find the degree of unsaturation

= 8 - 10/2   +1

= 8-5+1 = 4

so there is aromatic ring

there is quartet and triplet which indicates the the ethyl group is there

answer is ethyl benzene